A shoe shop has a sale.
X shoese Was $95
Sale price is 20% off
Y shose Was $90
Sale price is 25% off
What is the difference in the sale prices of these two pairs of shoes?

Answer by ikleyn(52788)

.
A shoe shop has a sale.
X shoes was $95
Sale price is 20% off
Y shoes was $90
Sale price is 25% off
What is the difference in the sale prices of these two pairs of shoes?
~~~~~~~~~~~~~~

Sale price for X shoes after discount is (1-0.2)*95 = 0.8*95 = 76 dollars.


Sale price for Y shoes after discount is (1-0.25)*90 = 0.75*95 = 67.50 dollars.


The difference in the sale prices of these two pairs of shoes is  76 - 67.50 = 8.50 dollars.


ANSWER.  The difference in the sale prices of these two pairs of shoes is  $8.50.

Solved.